A 4.5 x 107 kg meteor fell and struck the Earth near Flagstaff AZ, leaving a 180 m deep hole. If the meteor was moving at 20,000 m/s upon impact, with what average force did the meteor hit Earth?


1
Expert's answer
2021-11-30T09:49:48-0500

We can find the average force of the meteor when it hits the Earth from the work-kinetic energy theorem:


"KE_f-KE_i=W=-Fd,""0-\dfrac{1}{2}mv^2=-Fd,""F=\dfrac{mv^2}{2d}=\dfrac{4.5\times10^7\ kg\times(20000\ \dfrac{m}{s})^2}{2\times180\ m}=5\times10^{13}\ N."